OpenStudy (anonymous):

The monthly cost of driving a car depends on the number of miles driven. Lynn found that in May her driving cost was $380 for 300 mi and in June her cost was $500 for 900 mi. Assume that there is a linear relationship between the monthly cost C of driving a car and the distance driven d. (a) Find a linear equation that relates C and d. (b) Use part (a) to predict the cost of driving 1600 mi per month.

OpenStudy (jack1):

linear relationship therefore eqn of line is y = mx + b y = cost = C x = miles = d find m find b m = gradient = y2-y1 / x2 - x1

OpenStudy (anonymous):

m= 5

OpenStudy (jack1):

cool, so b = y intercept so use m = 5 to find b

OpenStudy (jack1):

use example y = 380, d = 300

OpenStudy (anonymous):

what?

OpenStudy (jack1):

y = mx + b $380 = 5 * 300miles + b 380 = 1500 + b b = 380 / 1500 = 0.2533

OpenStudy (anonymous):

so y=5(x)+0.2533

OpenStudy (jack1):

no, sorry, as m = 1/5 = (0.2) lets try again with that and get b = ...?

OpenStudy (jack1):

380 / 60 = b = 6.333

OpenStudy (anonymous):

how does m=1/5?

OpenStudy (jack1):

y2 - y1 / x2 -x1 = 500 - 380 / 900 - 300 = 120 / 600 = 1/5

OpenStudy (jack1):

as cost ( y ) or ( C ) is dependant on distance (x) or (d) whichever you prefer to call it

OpenStudy (jack1):

are u following ok with that? @hannahmarie620

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so c= 1/5(d)+6.33

OpenStudy (jack1):

perfect

OpenStudy (jack1):

now just plug in d as 1600 and ur set

OpenStudy (anonymous):

its wrong

OpenStudy (jack1):

i agree

OpenStudy (jack1):

i got 326 it looks soooo wrong, im sorry

OpenStudy (jack1):

aw dammit it's b = 380 MINUS 60, not divide by

OpenStudy (jack1):

so b = 320, sot 6.3

OpenStudy (jack1):

y = mx +b C = 1/5 d + b using $500 and 900 miles 500 = 1/5 (900) + b b = 500 - 180 = 320 so d = 1600 C = md + b C = 1/5 * 1600 + 320 C = 320 + 320 = $640

OpenStudy (jack1):

sorry for the screw up before i'll go to bed now

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

whats the equation

OpenStudy (anonymous):

nvm